# Lifelong Personalized Low-Rank Adaptation of Large Language Models for Recommendation

### v1

@@ Abstract.
We primarily focus on the field of large language models (LLMs) for recommendation, which has been actively explored recently and poses a significant challenge in effectively enhancing recommender systems with logical reasoning abilities and open-world knowledge. Current mainstream efforts mainly center around injecting personalized information from recommendation models into LLMs by customizing input templates or aligning representations between semantic and recommendation spaces at the prediction layer. However, they face three significant limitations: (1) LoRA is mostly used as a core component in existing works, but personalization is not well established in LoRA parameters as the LoRA matrix shared by every user may not cater to different users’ characteristics, leading to suboptimal performance. (2) Although lifelong personalized behavior sequences are ideal for personalization, their use raises effectiveness and efficiency issues since LLMs require escalating training and inference time to extend text lengths. (3) Existing approaches aren’t scalable for large datasets due to training efficiency constraints. Thus, LLMs only see a small fraction of the datasets (e.g., less than 10%) instead of the whole datasets, limiting their exposure to the full training space. To address these problems, we propose RecLoRA. This model incorporates a Personalized LoRA module that maintains independent LoRAs for different users and a Long-Short Modality Retriever that retrieves different history lengths for different modalities, significantly improving performance while adding minimal time cost. Furthermore, we design a Few2Many Learning Strategy, using a conventional recommendation model as a lens to magnify small training spaces to full spaces. Extensive experiments on public datasets demonstrate the efficacy of our RecLoRA compared to existing baseline models.
@@ 1. Introduction
Recommender systems play essential roles in various online services to mitigate the information overload problem and meet users’ information needs (Guo et al., 2017a; Lin et al., 2023b; Liu et al., 2024; Qin et al., 2022, 2023). Besides, large language models (LLMs) have experienced significant growth in the field of natural language processing (NLP), demonstrating a remarkable ability to comprehend human language and generate text that closely resembles human writings for a broad range of tasks (Zhao et al., 2023; Brown et al., 2020; Touvron et al., 2023). Recent studies have started to investigate the potential of LLMs for recommender systems under a variety of recommendation tasks, such as listwise ranking and pointwise scoring (Lin et al., 2023a, 2024b). These studies usually inject personalized knowledge from recommendation domains into LLMs. (Lin et al., 2024b; Zhang et al., 2023b; Zheng et al., 2023). This personalized knowledge plays a pivotal role in tailoring experiences to individual preferences and enhancing user interactions and engagements (Chen et al., 2023). So our crucial points are what useful personalized knowledge we can provide and how to inject it into LLM to help with predictions.
As the proverb ”Actions speak louder than words” suggests, a user’s behavior sequence is the most direct expression of their preferences. It typically provides the most detailed personalized information about users in recommendation scenarios. As illustrated in Figure 1, if a user’s behavior can be collected and modeled from birth to the present, their interests can be accurately predicted. Several mainstream studies have designed memory mechanisms to model lifelong behavior sequences (Ren et al., 2019; Pi et al., 2019). Other approaches employ retrieval methods to extract relevant information from entire lifelong sequences (Qin et al., 2020; Pi et al., 2020a; Chang et al., 2023). These methods prove the effectiveness of user behavior sequences. However, they fail to leverage general knowledge of LLMs to help users’ predictions.
@@
Firstly, personalization is not well established when adapting LLMs to recommender systems. For the sake of memory efficiency, mainstream approaches typically adopt low-rank adaptation (LoRA) techniques (Hu et al., 2021; Dettmers et al., 2023) and perform parameter-efficient finetuning (PEFT) to inject personalized knowledge from recommendation domains into LLMs (Lin et al., 2024b; Zhang et al., 2023b; Zheng et al., 2023). Ideally, each user should have a unique LoRA matrix to achieve personalized recommendations. However, current methods only implement prompt personalization or representation personalization, while the LoRA module, a core component of fine-tuning LLMs, remains static for all users and target items. This lack of parameter personalization results in inferior capabilities to effectively manage user dynamics and preference shifts in modern recommender systems.
Secondly, utilizing lifelong personalized behavior sequences in the prompt or representation layer raises issues of effectiveness and efficiency. Since LLMs are sensitive to input length and have a limit on input size, they tend to perform poorly on long behavior sequences, resulting in a ceiling on effectiveness.(Lin et al., 2024b) Additionally, as the length of the input behavior sequence and the volume of training data increase, the time required for training and inference escalates rapidly, leading to efficiency problems.
@@
To address the aforementioned issues, we propose a novel Personalized Low-rank Adaptation Framework of Large Language Models for Recommendation (i.e., RecLoRA). Specifically, we maintain a set of parallel, independent LoRA weights instead of a single static LoRA module as suggested in previous works (Lin et al., 2024b; Zhang et al., 2023b). As shown in Figure 1, this personalized LoRA module functions to incorporate personalized user knowledge into large language models, serving as a strategy for parameter personalization. We assign a sequential CRM and adapt the sequence representation to personalized meta-LoRA weights. More precisely, we use a soft routing method to aggregate meta-LoRA weights guided by the CRM. Instead of extending user behavior sequences in the text input, RecLoRA assigns longer sequences in the sequential CRM, which significantly improves effectiveness while adding minimal time cost. Moreover, the CRM is pre-trained on the full training space, so transforming the CRM to meta-LoRA weights acts like a lens, expanding the small fraction of training data exposed to the LLM to encompass the entire training space without increasing LLM training time.
@@
  • To the best of our knowledge, we are the first to consider LoRA parameters personalization in the recommendation task. We design a personalized LoRA module that includes a CRM on full training space and a soft routing method to aggregate meta-LoRA weights with CRM instructions.
  • We propose a novel framework, RecLoRA, which realizes a better trade-off between effectiveness and efficiency. RecLoRA has a Few2Many Magnifier where only few-shot training data is needed to efficiently tune LLM and a Long-short Modality Retriever module, which extends the sequence length on the ID input side to improve performance while almost not increasing the time cost.
  • Our personalized LoRA module is an easy-to-plugin module, which can be used in any LLM4rec model with LoRA modules.
We have conducted offline experiments on three publicly available datasets and achieved promising results, demonstrating the effectiveness of our proposed model.
@@ 2.1. Problem Formulation
The label $y_{i}\in\{1,0\}$ indicating the interaction signal of the user towards the item (i.e., prefer or not).
@@
According to different input data transformations, there generally exist two different recommendation paradigms: (1) The input of ID modality $x_{i}^{ID}$ obtained by one-hot encoding for conventional recommendation models (CRMs), and (2) The input of textual modality $x_{i}^{text}$ generated by hard prompt template for LLMs as recommenders.
@@ 2.3. Large Language Models as Recommenders
Typically, large language models (LLMs) refer to Transformer-based language models with at least billion-level parameters that are trained on massive text datasets, demonstrating remarkable capacity in various natural language tasks. When adapting directly LLMs as the recommenders, we need to convert the raw data $(x_{i},y_{i})$ into textual input-output pairs $(x_{i}^{text},y_{i}^{text})$ with hard prompt template. We illustrate one template example in Figure 2.
@@ 2.4. Low-Rank Adaptation of LLMs
Low-rank adaptation (LoRA) (Hu et al., 2021) serves as a popular parameter-efficient finetuning (PEFT) method to reduce the resource consumption of finetuning LLMs that possess massive parameters. The basic idea of LoRA is to maintain two trainable lightweight matrices $A$ and $B$ that are attached to a frozen pretrained weight matrix $W$. Hence, the linear transformation $Y=XW$ is reformulated as:
@@
so as to ensure the initial output $Y^{\prime}=XW+XAB^{T}$ equal to $Y=XW$. During finetuning, $W$ is fixed while $A$ and $B$ are updated by SGD-based optimization methods like Adam (Kingma and Ba, 2014).
When applying LoRA to LLM finetuning, we can selectively attach LoRA weights to target matrices of certain types inside the LLM. e.g., the matrices of query, key and value in the self-attention modules or the linear matrices in feed-forward networks (FFNs).
@@ 3. Methodology
In this section, we introduce our proposed RecLoRA (Personalized Low-rank Adaptation for Recommendation) framework and its training process in details.
@@ 3.1. Overview of RecLoRA
We propose the RecLoRA framework to address three major limitations: (1) To construct personalized low-rank adaptation for LLMs in recommendation, we employ a parallel meta-LoRA mechanism incorporating personalized knowledge. (2) To alleviate the effectiveness and efficiency issues associated with behavior sequence extension, we design the Long-Short Modality Retriever. (3) To expand the receptive field of LLMs to the entire training space, we develop the Few2Many Magnifier Strategy. In this section, we provide a comprehensive explanation of our RecLoRA framework, including its network architecture and training process.
Figure 3 illustrates the overall framework of the RecLoRA model. As is shown in the figure, RecLoRA takes two input components: samples in ID modality $x^{ID}$ and samples in text modality $x^{text}$. Each modality $x$ has user profile, candidate item, context features and user history behaviors as formulated in Eq. (1). So we start by transforming the recommendation dataset into these two modalities and then feed $x^{ID}$ and $x^{text}$ into CRM and LLM models respectively. During transformation, a Long-Short Modality Retriever retrieves long history for $x^{ID}$ and short history for $x^{test}$.
(7) $x^{ID}$ $=LSMR(u^{ID},v^{ID})$
$x^{text}$ $=LSMR(u^{text},v^{text})$
$N^{ID}_{u_{i}}$ $>N^{text}_{u_{i}}$
where $N$ is the sequence length and they would be described in detail in section 3.4.
@@
(9) $h_{j+1}$ $=Layer_{j}(h_{j})+LoRA(h_{j})$
$h_{j+1}$ $=Layer_{j}(h_{j})+PLoRA(h_{j},x^{ID})$
@@ 3.2. Personalized Low-Rank Adaption
However, this approach is non-scalable, as its model complexity grows linearly with the number of users, leading to a tremendous amount of memory consumption. Moreover, it overlooks the overlap and commonality among behavior patterns of different users, which can lead to suboptimal performance.
@@
Then, we can obtain the personalized LoRA weights by combining them with a trainable gating network.
(11) $A_{p}B_{p}^{T}=\sum_{k=1}^{N_{m}}\alpha_{k}~{}A_{k}B_{k}^{T}.$
where $A_{p}B_{p}$ are the final personalized LoRA matrix, and $N_{m}$ is the number of meta-LoRA weights.
In this equation, the $k_{th}$ gating weight $\alpha_{k}$ represents the importance and relevance of the current user and the $k_{th}$ LoRA weight, acting as an indicator for personalization. Thus, To capture the complicated correlations and achieve personalized aggregation of the meta-LoRA weights, we introduce the conventional recommendation model (CRM) to provide personalized representations for the gating network. CRM takes discrete ID $x^{ID}$ as input and generates hidden states as personalized representations.
(12) $R_{c}=CRM(x^{ID})$
where $R_{c}\in\mathbb{R}^{n\times d_{c}}$ represents the CRM’s output. Specifically, we use a sequential behavior model, SIM(Pi et al., 2020b), as user history provides the most ideal personalization, and SIM is a state-of-the-art model for sequential behavior in recommendations.
After obtaining the representation, we adapt it to the gating weight to guide the meta-LoRA weights with personalized knowledge. We use a feed-forward network (FFN) with ReLU activation and LayerNorm as the adapter:
(13) $\beta$ $=W_{2}\times\text{ReLU}(\text{LN}(W_{1}R_{c}+b_{1})))+b_{2}$
@@
(14) $\alpha_{j}$ $=\frac{\exp(\frac{\beta_{j}}{tp})}{\sum_{j^{\prime}=1}^{N_{m}}\exp(\frac{\beta_{j^{\prime}}}{tp})}$
where $W_{1}\in\mathbb{R}^{d_{c}\times d_{h}}$, $W_{2}\in\mathbb{R}^{d_{h}\times d_{r}}$, $b_{1}\in\mathbb{R}^{d_{h}}$, $b_{2}\in\mathbb{R}^{d_{r}}$, and $\alpha\in\mathbb{R}^{n\times N_{m}}$ is the gating weight of Eq. (11). $tp$ is a temperature hyperparameter in the softmax function, making it smoother or sharper. Most importantly, softmax guarantees the distribution of personalized LoRA weights matches each meta-LoRA weights.
@@ 3.3. Few2Many Magnifier Strategy
In the first stage, we train a conventional recommendation model using the full training dataset to obtain a well-trained model capable of providing sample-level personalized vector representations with high generalization based on the input ID modal data $x^{ID}$. This is achieved through a binary cross-entropy loss
(15) $\mathcal{L}^{ID}=\sum_{x_{i}^{ID}\in\mathcal{D}}[-y_{i}^{ID}\cdot\log(\sigma(\hat{y}_{i}^{ID}))-(1-y_{i}^{ID})\cdot\log(1-\sigma(\hat{y}_{i}^{ID}))].$
where $\sigma$ is sigmoid function, $y_{i}^{ID}$ is the true label and $\hat{y}_{i}^{ID}$ is the prediction label calculated in Eq. (3) Now CRM has fitted the full training space and learned most of the knowledge contained in the whole dataset.
In the second stage, we downsample the large-scale training set to obtain a smaller training set for the efficient parameter tuning of RecLoRA, which includes large language models. During this process, the fully trained traditional recommendation model from the first stage provides personalized information, ensuring full spatial perspectives. This information is combined to form a personalized LoRA matrix with sufficient generalization, as detailed in Section 3.2. Thus, even though the large language model sees only a small number of training samples during fine-tuning, it extends its receptive field to the full training space through the conventional recommendation models and personalized LoRA matrices. This greatly enhances the sample efficiency and recommendation performance of the large language model. It should be noted that during this process, only the personalized LoRA module is updated, while the traditional recommendation models and large language models remain fixed. The training loss used is the cross-entropy loss in traditional causal language modeling.
During the inference stage, large language model doesn’t directly give a pointwise score $\hat{y}^{text}\in\{0,1\}$. Therefore, we intercept the vocabulary scores, and then conduct a bidimensional softmax over the scores of binary key answer words. Specifically, the scores for “Yes” and “No” are $s_{y}$ and $s_{n}$ respectively. Then the pointwise scoring of LLMs can be written as:

### v2

@@ Abstract.
We primarily focus on the field of large language models (LLMs) for recommendation, which has been actively explored recently and poses a significant challenge in effectively enhancing recommender systems with logical reasoning abilities and open-world knowledge. Current mainstream efforts mainly center around injecting personalized information from recommendation models into LLMs by customizing input templates or aligning representations between semantic and recommendation spaces at the prediction layer. However, they face three significant limitations: (1) LoRA is mostly used as a core component in existing works, but personalization is not well established in LoRA parameters as the LoRA matrix shared by every user may not cater to different users’ characteristics, leading to suboptimal performance. (2) Although lifelong personalized behavior sequences are ideal for personalization, their use raises effectiveness and efficiency issues since LLMs require escalating training and inference time to extend text lengths. (3) Existing approaches fail to be scalable for large datasets due to training efficiency constraints. Thus, LLMs only see a small fraction of the datasets (e.g., less than 10%) instead of the whole datasets, limiting their exposure to the full training space. To address these problems, we propose RecLoRA. This model incorporates a personalized LoRA module that maintains independent LoRAs for different users and a Long-Short Modality Retriever that retrieves different history lengths for different modalities, significantly improving performance with negligible time cost increase. Furthermore, we design a Few2Many Learning Strategy, using a conventional recommendation model as a lens to magnify small training spaces to full spaces. Extensive experiments on real-world datasets demonstrate the efficacy of our RecLoRA compared to strong baseline models.
@@ 1. Introduction
Recommender systems play essential roles in various online services to mitigate the information overload problem and meet users’ information needs (Guo et al., 2017a; Lin et al., 2023b; Liu et al., 2024; Qin et al., 2022, 2023). Besides, large language models (LLMs) have experienced significant growth in the field of natural language processing (NLP), demonstrating a remarkable ability to comprehend human languages and generate texts that closely resemble human writings for a broad range of tasks (Zhao et al., 2023; Brown et al., 2020; Touvron et al., 2023). Recent studies have started to investigate the potential of LLMs for recommender systems under various recommendation tasks, such as listwise ranking and pointwise scoring (Lin et al., 2023a, 2024b). These studies usually inject personalized knowledge from recommendation domains into LLMs (Lin et al., 2024b; Zhang et al., 2023b; Zheng et al., 2023). This personalized knowledge plays a pivotal role in tailoring experiences to individual preferences and enhancing user interactions and engagements (Chen et al., 2023). So our crucial points are what useful personalized knowledge we can provide and how to inject it into LLM to help with predictions.
As the proverb ”Actions speak louder than words” suggests, a user’s behavior sequence is the most direct expression of their preferences. It typically provides the most detailed personalized information about users in recommendation scenarios. As illustrated in Figure 1, if a user’s behaviors can be collected and modeled from birth to the present, their interests can be accurately predicted. Several mainstream studies have designed memory mechanisms to model lifelong behavior sequences (Ren et al., 2019; Pi et al., 2019). Other approaches employ retrieval methods to extract relevant information from entire lifelong sequences (Qin et al., 2020; Pi et al., 2020a; Chang et al., 2023). These methods prove the effectiveness of user behavior sequences, which also shows potential in LLM-based recommendation models.
@@
Firstly, personalization is not well established when adapting LLMs to recommender systems. For the sake of memory efficiency, mainstream approaches typically adopt low-rank adaptation (LoRA) techniques (Hu et al., 2021; Dettmers et al., 2023) and perform parameter-efficient finetuning (PEFT) to inject personalized knowledge from recommendation domains into LLMs (Lin et al., 2024b; Zhang et al., 2023b; Zheng et al., 2023). In the context of PEFT, maintaining a unique LoRA matrix for each user can enable maximum personalization, effectively capturing and adapting to unique user features and changing interests. However, current methods only implement prompt personalization or representation personalization, while the LoRA module, a core component of fine-tuning LLMs, remains static for all users and target items. This lack of parameter personalization results in inferior capabilities to effectively manage user dynamics and preference shifts in modern recommender systems.
Secondly, utilizing lifelong personalized behavior sequences in the prompt or representation layer raises issues of effectiveness and efficiency(Lin et al., 2024b). Since LLMs are sensitive to input length and have a limit on input size, they tend to perform poorly on long behavior sequences, resulting in a ceiling on effectiveness. Additionally, as the length of the input behavior sequence increases, the time required for training and inference escalates rapidly, leading to efficiency problems.
@@
To address the aforementioned issues, we propose a novel Personalized Low-rank Adaptation Framework of Large Language Models for Recommendation (i.e., RecLoRA). Specifically, we maintain a set of parallel, independent LoRA weights instead of a single static LoRA module as suggested in previous works (Lin et al., 2024b; Zhang et al., 2023b). As shown in Figure 1, this personalized LoRA module functions to incorporate personalized user knowledge into large language models, serving as a strategy for parameter personalization. We assign a sequential CRM and adapt the user representation to personalized LoRA weights. More precisely, we use a soft routing method to aggregate meta-LoRA weights guided by the CRM. Instead of extending user behavior sequences in the input prompt, RecLoRA assigns longer sequences in the sequential CRM, significantly improving effectiveness with negligible cost increase. Moreover, the CRM is pre-trained on the full training space, so dynamically generating LoRA weights based on the CRM can act like a magnifier, enlarging the training signal observed from a small fraction of data to the entire data space without increasing LLM training time.
@@
  • To the best of our knowledge, we are the first to consider LoRA parameters personalization in the recommendation task. We design a personalized LoRA module that includes a CRM trained on full training space and a soft routing method to aggregate meta-LoRA weights with CRM instructions.
  • We propose a novel framework, RecLoRA, which realizes a better trade-off between effectiveness and efficiency. RecLoRA has a Few2Many Magnifier where only few-shot training data is needed to efficiently tune LLM and a Long-short Modality Retriever module, which extends the sequence length in the CRM to improve performance with negligible time cost increase.
  • Our personalized LoRA module is an easy-to-plugin module, which can be used in any LLM-based recommendation model with LoRA modules.
Extensive experiments on three public real-world datasets have been conducted, where RecLoRA achieves consistent superiority over strong baselines.
@@ 2. Preliminary
In this section, we will formulate the problem and introduce Low-Rank Adaptation, which is the basis of RecLoRA.
@@ 2.1. Problem Formulation
The label $y_{i}\in\{1,0\}$ indicates the interaction signal of the user towards the item (i.e., prefer or not).
@@
According to different input data transformations, there generally exist two different recommendation paradigms: (1) The input of ID modality $x_{i}^{ID}$ obtained by one-hot encoding for CRMs (2) The input of textual modality $x_{i}^{text}$ generated by hard prompt template for LLMs as recommenders.
@@ 2.3. Large Language Models as Recommenders
Typically, large language models (LLMs) refer to Transformer-based language models with at least billion-level parameters that are trained on massive text datasets, demonstrating a remarkable capacity in various natural language tasks. When adopting directly LLMs as the recommenders, we need to convert the raw data $(x_{i},y_{i})$ into textual input-output pairs $(x_{i}^{text},y_{i}^{text})$ with a hard prompt template. We illustrate one template example in Figure 2.
@@ 2.4. Low-Rank Adaptation of LLMs
Low-rank adaptation (LoRA) (Hu et al., 2021) serves as a popular parameter-efficient finetuning (PEFT) method to reduce the resource consumption of finetuning LLMs that possess massive parameters. The basic idea of LoRA is to maintain two trainable lightweight matrices $A$ and $B$ that are attached to a frozen pre-trained weight matrix $W$. Hence, the linear transformation $Y=XW$ is reformulated as:
@@
so as to ensure the initial output $Y^{\prime}=XW+XAB^{T}$ equals to $Y=XW$. During fine-tuning, $W$ is fixed while $A$ and $B$ are updated by SGD-based optimization methods like Adam (Kingma and Ba, 2014).
When applying LoRA to LLM finetuning, we can selectively attach LoRA weights to target matrices of certain types inside the LLM. e.g., the matrices of query, key, and value in the self-attention modules or the linear matrices in feed-forward networks (FFNs).
@@ 3. Methodology
In this section, we introduce our proposed RecLoRA (Personalized Low-rank Adaptation for Recommendation) framework and its training process in detail.
@@ 3.1. Overview of RecLoRA
We propose the RecLoRA framework to address three major limitations: (1) To construct personalized low-rank adaptation for LLMs in recommendation, we employ a parallel meta-LoRA mechanism incorporating personalized knowledge. (2) To alleviate the effectiveness and efficiency issues associated with behavior sequence extension, we design a Long-Short Modality Retriever. (3) To expand the receptive field of LLMs to the entire training space, we develop a Few2Many Magnifier Strategy. In this section, we provide a comprehensive explanation of our RecLoRA framework, including its network architecture and training process.
Figure 3 illustrates the overall framework of the RecLoRA model. As is shown in the figure, RecLoRA takes two input components: samples in ID modality $x^{ID}$ and samples in text modality $x^{text}$. Each modality $x$ has a user profile, a candidate item, context features, and user history behaviors as formulated in Eq. (1). So we start by transforming the recommendation dataset into these two modalities and then feed $x^{ID}$ and $x^{text}$ into CRM and LLM models respectively. During transformation, a Long-Short Modality Retriever retrieves long history for $x^{ID}$ and short history for $x^{test}$.
(7) $x^{ID}$ $=LSMR(u^{ID},v^{ID}),$
$x^{text}$ $=LSMR(u^{text},v^{text}),$
$N^{ID}_{u_{i}}$ $>N^{text}_{u_{i}},$
where $N$ is the sequence length and this module would be described in detail in section 3.4.
@@
(9) $h_{j+1}$ $=Layer_{j}(h_{j})+LoRA(h_{j}),$
$h_{j+1}$ $=Layer_{j}(h_{j})+PLoRA(h_{j},x^{ID}),$
@@ 3.2. Personalized Low-Rank Adaption
However, this approach is non-scalable, as its model complexity grows linearly with the number of users, leading to a tremendous amount of memory consumption. Moreover, it overlooks the overlap and commonality among behavior patterns of different users, which can lead to suboptimal performances.
@@
Then, we can obtain the personalized LoRA weights by combining them with a trainable gating network as
(11) $A_{p}B_{p}^{T}=\sum_{k=1}^{N_{m}}\alpha_{k}~{}A_{k}B_{k}^{T},$
where $A_{p},B_{p}$ are the final personalized LoRA matrices, and $N_{m}$ is the number of meta-LoRA weights.
In this equation, the $k_{th}$ gating weight $\alpha_{k}$ represents the importance and relevance of the current user and the $k_{th}$ LoRA weight, acting as an indicator for personalization. Thus, to capture the complicated correlations and achieve personalized aggregation of the meta-LoRA weights, we introduce the conventional recommendation model (CRM) to provide personalized representations for the gating network. CRM takes inputs of discrete ID and generates hidden states as personalized representations.
(12) $R_{c}=CRM(x^{ID}),$
where $R_{c}\in\mathbb{R}^{n\times d_{c}}$ represents the CRM’s output. Specifically, we use a sequential behavior model, SIM(Pi et al., 2020b), as user history provides the most ideal personalization, and SIM is a state-of-the-art model for sequential behaviors in recommendations.
After obtaining the representation, we adapt it to the gating weight to guide the meta-LoRA weights with personalized knowledge. We use a feed-forward network (FFN) with a ReLU activation and a LayerNorm as the adapter:
(13) $\beta$ $=W_{2}\times\text{ReLU}(\text{LN}(W_{1}R_{c}+b_{1})))+b_{2},$
@@
(14) $\alpha_{j}$ $=\frac{\exp(\frac{\beta_{j}}{tp})}{\sum_{j^{\prime}=1}^{N_{m}}\exp(\frac{\beta_{j^{\prime}}}{tp})},$
where $W_{1}\in\mathbb{R}^{d_{c}\times d_{h}}$, $W_{2}\in\mathbb{R}^{d_{h}\times d_{r}}$, $b_{1}\in\mathbb{R}^{d_{h}}$, $b_{2}\in\mathbb{R}^{d_{r}}$, and $\alpha\in\mathbb{R}^{n\times N_{m}}$ is the gating weight of Eq. (11). $tp$ is a temperature hyperparameter in the softmax function, making it smoother or sharper. Most importantly, softmax guarantees the distribution of personalized LoRA weights matches each meta-LoRA weight.
@@ 3.3. Few2Many Magnifier Strategy
In the first stage, we train a conventional recommendation model using the full training dataset to obtain a well-trained model capable of providing sample-level personalized vector representations with high generalization based on the inputs $x^{ID}$ in ID modality . This is achieved through a binary cross-entropy loss
(15) $\mathcal{L}^{ID}=\sum_{x_{i}^{ID}\in\mathcal{D}}[-y_{i}^{ID}\cdot\log(\sigma(\hat{y}_{i}^{ID}))-(1-y_{i}^{ID})\cdot\log(1-\sigma(\hat{y}_{i}^{ID}))],$
where $\sigma$ is the sigmoid function, $y_{i}^{ID}$ is the true label and $\hat{y}_{i}^{ID}$ is the prediction label calculated in Eq. (3) Now, the CRM has fit the full training space and has learned most of the knowledge contained in the whole dataset.
In the second stage, we downsample the large-scale training set to obtain a smaller training set for the efficient parameter tuning of RecLoRA, which includes large language models. During this process, the fully trained traditional recommendation model from the first stage provides personalized information, ensuring full spatial perspectives. This information is combined to form a personalized LoRA matrix with sufficient generalization, as detailedly described in Section 3.2. Thus, even though the large language model sees only a small number of training samples during fine-tuning, it extends its receptive field to the full training space through the conventional recommendation models and personalized LoRA matrices. This greatly enhances the sample efficiency and recommendation performance of the large language model. It should be noted that during this process, only the personalized LoRA module is updated, while the traditional recommendation models and large language models remain fixed. The training loss used is the cross-entropy loss in traditional causal language modeling.
During the inference stage, the large language model does not directly give a pointwise score $\hat{y}^{text}\in\{0,1\}$. Therefore, we intercept the vocabulary scores, and then conduct a bidimensional softmax over the scores of binary key answer words. Specifically, the scores for “Yes” and “No” are $s_{y}$ and $s_{n}$ respectively. Then, the pointwise scoring of LLMs can be written as: